# Hybrid Beamforming Design for Near-Field ISAC with Modular XL-MIMO

### v1

@@ IV-D SDR-based Randomization for Near-Optimal Solution
$\ \alpha_{1}^{2}\text{tr}\left({{\mathbf{R}}_{\text{BB}}}{{\mathbf{\Phi}}_{1}}\right)-{{\Gamma}_{s}}\sum\limits_{q=2}^{Q}{\alpha_{q}^{2}\text{tr}\left({{\mathbf{R}}_{\text{BB}}}{{\mathbf{\Phi}}_{q}}\right)}\geq{{\Gamma}_{0}},$ (59c)
@@
The relaxed problem (59) is convex and can be efficiently solved using interior-point method. However, the optimal solution $\mathbf{R}_{\text{BB}}^{*}$ to the relaxed problem may not satisfy the rank constraint $\mathrm{rank}(\mathbf{R}_{\text{BB}})=N_{s}$. To recover a rank-constrained solution to the original problem (40), we apply the randomization technique, which generates a set of candidate solutions from $\mathbf{R}_{\text{BB}}^{*}$ and selects the one that maximizes the objective function (59a) while satisfying the constraints (59b) and (59c). The detailed procedure of the proposed algorithm, referred to as Low-Complexity SDR-based Randomization for Rank-constrained Solution (LC-SDR-RRS), is summarized in Algorithm 2.
Algorithm 2 LC-SDR-RRS Algorithm
0:  Problem data: $\mathbf{H}_{c}$, $\mathbf{\tilde{U}}$, $\sigma_{c}^{2}$, $N_{s}$, $M$, $\Gamma_{s}$, $\alpha_{q}$, $\mathbf{\Phi}_{q}$, $\sigma_{s}^{2}$, $\mathbf{w}$.
0:  Approximate solution $\mathbf{W}_{\text{BB}}$ to problem (40) with $\mathrm{rank}(\mathbf{W}_{\text{BB}}\mathbf{W}_{\text{BB}}^{H})=N_{s}$.
1:  Solve the SDP problem (59) using interior-point method to obtain $\mathbf{R}_{\text{BB}}^{*}$;
2:  Compute the eigenvalue decomposition of $\mathbf{R}_{\text{BB}}^{*}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{H}$;
3:  Let $\mathbf{U}_{N_{s}}$ be the matrix containing the eigenvectors corresponding to the $N_{s}$ largest eigenvalues;
4:  Set $\mathbf{V}=\mathbf{U}_{N_{s}}\mathbf{\Lambda}_{N_{s}}^{1/2}$, where $\mathbf{\Lambda}_{N_{s}}$ is the diagonal matrix containing the $N_{s}$ largest eigenvalues;
5:  for $i=1,\dots,10N_{s}$ do
6:     Generate ai random matrix $\mathbf{Z}_{i}\in\mathbb{C}^{N_{s}\times N_{s}}$ with i.i.d. entries drawn from $\mathcal{CN}(0,1)$;
7:     Set $\mathbf{W}_{i}=\mathbf{V}\mathbf{Z}_{i}$;
8:     Scale $\mathbf{W}_{i}$ to satisfy the power constraint (59b):$\mathbf{W}_{i}\leftarrow\sqrt{\frac{N_{s}}{M\cdot\textit{tr}(\mathbf{W}_{i}\mathbf{W}_{i}^{H})}}\mathbf{W}_{i}$;
9:  end for
10:  Choose the best solution among $\{\mathbf{W}_{i}\}_{i=1}^{10N_{s}}$ that satisfies the SCNR constraint (59c) and maximizes the objective (40a):
@@
11:  return  $\mathbf{W}_{\text{BB}}=\mathbf{W}^{*}$
@@
Compared to the manifold optimization approach introduced in subsection IV-C, the proposed LC-SDR-RRS algorithm has several advantages. First, by relaxing the non-convex rank constraint, the LC-SDR-RRS transforms the original problem into a convex SDP problem, which can be globally solved in polynomial time. Second, the randomization technique employed in the LC-SDR-RRS algorithm enables the generation of multiple candidate solutions, enhancing the probability of finding a high-quality solution to the original problem. Finally, the LC-SDR-RRS exhibits reduced sensitivity to initialization and can provide a favorable starting point for the manifold optimization approach, potentially accelerating its convergence and improving the solution quality.
@@ V Simmulation Results
In this section, we extensively evaluate the communication and sensing performance of the modular XL-MIMO ISAC system with the proposed algorithms, and investigate the impact of various system parameters. We consider a downlink sensing scenario, where the BS communicates with a multi-antenna user while sensing a target in the presence of interferences. Unless otherwise specified, the BS located at the origin is equipped with a modular XL-array consisting of $K=4$ subarrays, and the number of antennas within each subarray is $M=32$. The central frequency is $38$ GHz, and the inter-antenna spacing is $d=\lambda/2\approx 0.00395$ m. The user is equipped with $N_{c}=16$ antennas and located at a distance of 40 m and an angle of $15^{\circ}$ relative to the Tx center. For the communication channel, we consider one LoS path and three NLoS paths. The scatterers are randomly distributed within a distance range of 5 to 30 m and an angle range of -60° to 60°. The path gains are generated according to the 3GPP TR 38.901 specification[30]. The target of interest is at the direction of $30^{\circ}$ with the distance of 30 m, and there exist two interferences at the same range as the target, with angles of 40° and -30°, respectively. The noise power for communication and sensing are set to -30 dBm and -20 dBm, respectively. For simplicity, we adopt the omnidirectional transmission with $\mathbf{R}_{X}=\mathbf{I}$ to calculate the fixed receive beamformer $\mathbf{w}$ according to (21) in the simulations. Once the optimal transmit covariance matrix $\mathbf{R}_{X}^{*}$ is obtained, we substitute it into (21) to derive the corresponding optimal receive beamformer $\mathbf{w}^{*}$.
@@
Fig. 2 illustrates the convergence behavior of Algorithm 2 under various barrier parameters and initial points. Fig. 2 (a) demonstrates the impact of the barrier parameter $t$ on the convergence. It is observed that a smaller $t$ leads to faster convergence but may result in a higher final cost, while a larger $t$ slows down the convergence but yields a lower final cost. This trade-off between convergence speed and solution quality is consistent with the theory of interior-point methods. The results suggest that a well-tuned barrier parameter, e.g., $t=1$, can strike a balance between convergence speed and optimality. Additionally, the algorithm exhibits stable convergence for a wide range of $t$ values from 0.5 to 100, indicating its robustness to the choice of barrier parameter. Fig. 2 (b) reveals that the proposed algorithm converges within limited iterations for all considered initial points, showcasing its robustness to initialization. However, the choice of initial point does affect the convergence speed and the final objective value. To improve the solution quality, we run the algorithm with different random initial points and select the best local optimum.
@@
Then, we investigate the impact of different subarray distributions on the performance of ISAC systems. With a fixed number of subarrays $K$ and antennas per subarray $M$, we consider three modular array configurations: random subarray layout (denoted as ”random”), where the subarrays are randomly distributed; uniform subarray layout (denoted as ”uniform”), where the subarrays are uniformly distributed; and collocated subarray layout (denoted as ”collocated”), where the subarrays are closely spaced with a half-wavelength inter-subarray spacing. For notational convenience, we denote Algorithm 1 as RM-JGD and Algorithm 2 as SDR-RRS.
Fig. 3 illustrates the user’s communication spectral efficiency versus the received SNR for both SDR-RRS and RM-JGD under three modular array configurations. It is noted that for both algorithms, the modular arrays with random and uniform subarray layouts achieve similar communication spectral efficiencies, surpassing the collocated subarray layout. Moreover, the performance gap widens as the SNR increases, indicating that for modular arrays with the same $K$ and $M$, widely spaced subarray distributions outperform the traditional collocated distribution. This superiority stems from the more pronounced near-field spherical wave characteristics of widely spaced subarrays, which enhance the channel rank and spatial multiplexing gain. Furthermore, SDR-RRS consistently outperforms RM-JGD across all array configurations, showcasing its ability to better handle the non-convex optimization problem and find higher-quality locally optimal solutions.
@@
Fig. 5 illustrates the impact of subarray scale on the communication performance of SDR-RRS algorithm in modular XL-MIMO ISAC systems. The total array aperture $S$ and the number of antennas $N_{t}=192$ are kept constant, while the number of antennas per subarray $M$ is varied as 16, 24, 32, 48, and 64, resulting in the corresponding number of subarrays $K$ being 12, 8, 6, 4, and 3, respectively. It can be observed that the optimal subarray scale for maximizing the communication spectral efficiency varies with the user distance. For user distances less than $25\mathrm{m}$, the subarray scale of $M=48$ outperforms the others, while for distances beyond $25\mathrm{m}$, the subarray scale of $M=64$ achieves the highest communication spectral efficiency. A larger number of subarrays $K$ provides higher spatial multiplexing gains, but a smaller $M$ reduces the beamforming gain. Thus, the optimal subarray scale at different distances represents a trade-off between spatial multiplexing and beamforming gains for maximizing the communication performance. Moreover, the performance gap between subarray scales diminishes with increasing distance due to the transition from near-field spherical wavefronts, which enhance spatial multiplexing for larger $K$ at shorter ranges, to far-field planar wavefronts where such gains diminish.
Figure 6: Communication spectral efficiency versus user distance for different numbers of NLoS paths.
Fig. 6 illustrates the impact of user distance and the number of NLoS paths on the communication spectral efficiency for both the SDR-RRS and RM-JGD algorithms. The communication spectral efficiency of both Algorithms decreases with increasing user distance due to increased path loss and diminishing near-field spherical wave characteristics, which reduce the channel rank and limit the number of independent spatial streams available for communication. It is noted that SDR-RRS consistently outperforms RM-JGD for all considered scenarios, showcasing its superior performance. Moreover, the performance gap between the two algorithms widens as the number of NLoS paths increases, highlighting the ability of SDR-RRS to better exploit multipath propagation for higher communication spectral efficiencies. It also reveals that even for distant users, a rich scattering environment with abundant NLoS paths can sustain high communication spectral efficiencies, highlighting the potential of exploiting multipath propagation to enhance cell-edge performance.
Figure 7: Communication spectral efficiency versus sensing SCNR threshold for different levels of overlap between communication channel scatterers and sensing interference.
We then investigate the impact of the overlap between communication channel scatterers and sensing interference on the communication-sensing performance tradeoff. The overlap parameter represents the number of scatterers that coincide with the sensing interference, with overlap values 0 indicating no overlap and thus low correlation between the communication and sensing channels. Overlap values of 1 and 2 represent that one and two scatterers in the environment also act as interference sources for the sensing targets, respectively. As shown in Fig. 7, for both the SDR-RRS and RM-JGD algorithms, a higher overlap value yields a higher communication spectral efficiency at the same sensing SCNR threshold, indicating that the system achieves a better communication-sensing tradeoff. This is because that the correlation between the communication and sensing channels increase as the number of overlapping scatterers and interference sources grows, leading to more efficient resource utilization for the proposed ISAC system.
In Fig. 8, the normalized mutiple signal classification (MUSIC) spectrum obtained by the proposed SDR-RRS algorithm is compared for $M=32$ with varying numbers of subarrays $K$, over a grid spanning $x\in[0:0.06:30]$ m and $y\in[0:0.06:30]$ m. The peak of the spectrum is consistently observed at the actual target location $(20\mathrm{m},\pi/4)$ for all values of $K$. Notably, the main lobe width containing the peak value narrows as $K$ increases, indicating enhanced range resolution at the same angular direction. This behavior can be attributed to the spherical wavefront characteristic across subarrays, which becomes more pronounced with increasing $K$, leading to improved range estimation accuracy. The results demonstrate that augmenting the number of subarrays can significantly enhance the range resolution capability of the MUSIC algorithm, thereby enabling more precise target localization.
@@ VI Conclusion
In this paper, we have proposed a novel modular XL-MIMO ISAC framework and developed the low-complexity hybrid beamforming algorithms. By exploiting the structural similarity between the communication and sensing channels based on the HSPM, we have derived a closed-form solution for the optimal analog beamforming matrix and transformed the joint analog-digital design problem into a lower-dimensional digital beamformer optimization. To address the rank-constrained digital beamformer optimization, we have proposed a manifold-based algorithm for reduced complexity and a SDR-based method for near-optimal solutions.
Extensive simulations have validated the effectiveness of the proposed framework and algorithms. The simulations reveal that increasing the number of subarrays enhances the spatial multiplexing gain and improves the range resolution for sensing. The presence of NLoS paths and the correlation between communication and sensing channels are shown to have a significant impact on performances of the ISAC system. Moreover, the optimal subarray scale is found to vary with the user distance, highlighting the trade-off between beamforming gain and spatial multiplexing gain. These findings offer useful guidelines for the deployment of modular XL-MIMO ISAC systems in various practical scenarios.
Appendix A Proof of Theorem 1
Based on the method proposed in [31, App. C], let
$\mathbf{R}_{X}^{*}=\bm{\Delta}\bm{\Delta}^{\mathrm{H}}.$ (60)
We can decompose $\bm{\Delta}$ as

### v2

@@ IV-D SDR-based Randomization for Near-Optimal Solution
$\ \alpha_{1}^{2}\text{tr}\left({{\mathbf{R}}_{\text{BB}}}{{\mathbf{\Phi}}_{0}}\right)-{{\Gamma}_{s}}\sum\limits_{q=1}^{Q-1}{\alpha_{q}^{2}\text{tr}\left({{\mathbf{R}}_{\text{BB}}}{{\mathbf{\Phi}}_{q}}\right)}\geq{{\Gamma}_{0}},$ (54c)
@@
The relaxed problem (54) is convex and can be efficiently solved using interior-point method. However, the optimal solution $\mathbf{R}_{\text{BB}}^{*}$ to the relaxed problem may not satisfy the rank constraint $\mathrm{rank}(\mathbf{R}_{\text{BB}})=N_{s}$. To recover a rank-constrained solution to the original problem (35), we apply the randomization technique, which generates a set of candidate solutions from $\mathbf{R}_{\text{BB}}^{*}$ and selects the one that maximizes the objective function (54a) while satisfying the constraints (54b) and (54c). The detailed procedure of the proposed algorithm, referred to as SDR-based Randomization for Rank-constrained Solution (SDR-RRS), is summarized in Algorithm 2.
Algorithm 2 SDR-RRS Algorithm
1:  Initialize $\mathbf{H}_{c}$, $\mathbf{\tilde{U}}$, $\sigma_{c}^{2}$, $N_{s}$, $M$, $\Gamma_{s}$, $\alpha_{q}$, $\mathbf{\Phi}_{q}$, $\sigma_{s}^{2}$, $\mathbf{w}$.
2:  Solve the SDP problem (54) using interior-point method to obtain $\mathbf{R}_{\text{BB}}^{*}$;
3:  Compute the eigenvalue decomposition of $\mathbf{R}_{\text{BB}}^{*}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{H}$;
4:  Let $\mathbf{U}_{N_{s}}$ be the matrix containing the eigenvectors corresponding to the $N_{s}$ largest eigenvalues;
5:  Set $\mathbf{V}=\mathbf{U}_{N_{s}}\mathbf{\Lambda}_{N_{s}}^{1/2}$, where $\mathbf{\Lambda}_{N_{s}}$ is the diagonal matrix containing the $N_{s}$ largest eigenvalues;
6:  for $i=1,\dots,10N_{s}$ do
7:     Generate a random matrix $\mathbf{Z}_{i}\in\mathbb{C}^{N_{s}\times N_{s}}$ with i.i.d. entries drawn from $\mathcal{CN}(0,1)$;
8:     Set $\mathbf{W}_{i}=\mathbf{V}\mathbf{Z}_{i}$;
9:     Scale $\mathbf{W}_{i}$ to satisfy the power constraint (54b):$\mathbf{W}_{i}\leftarrow\sqrt{\frac{N_{s}}{M\cdot\textit{tr}(\mathbf{W}_{i}\mathbf{W}_{i}^{H})}}\mathbf{W}_{i}$;
10:  end for
11:  Choose the best solution among $\{\mathbf{W}_{i}\}_{i=1}^{10N_{s}}$ that satisfies the SCNR constraint (54c) and maximizes the objective (35a):
@@
12:  Output $\mathbf{W}_{\text{BB}}=\mathbf{W}^{*}$
@@
Compared to the RM-JGD algorithm, the proposed SDR-RRS algorithm has several advantages. First, by relaxing the non-convex rank constraint, the SDR-RRS transforms the original problem into a convex SDP problem, which can be globally solved in polynomial time. Second, the randomization technique in SDR-RRS generates multiple candidate solutions based on the global optimum of the relaxed problem, ensuring better solution than RM-JGD, where the obtained local optimum highly depends on random initialization and may be far from the global optimum.
@@ V Simmulation Results
In this section, we evaluate the performance of the proposed modular XL-MIMO ISAC system and algorithms through extensive simulations. Unless otherwise specified, the BS is equipped with a modular XL-array of $K=6$ subarrays, each containing $M=32$ antennas. The operating frequency is 38 GHz with half-wavelength antenna spacing $d=\lambda/2\approx 0.00395$ m. The user is equipped with $N_{c}=16$ antennas and located at a distance of 40 m and an angle of $15^{\circ}$ relative to the Tx center. The channel consists of one LoS and three NLoS paths, with scatterers randomly distributed between 5-30 m and $-60^{\circ}$ to $60^{\circ}$. The path gains follow the 3GPP TR 38.901 specification [32]. The target is positioned at 30 m and $30^{\circ}$, with two interferences at the same range but at angles of $40^{\circ}$ and $-30^{\circ}$. The noise power for communication and sensing are set to -30 dBm and -20 dBm, respectively. For simplicity, we adopt the omnidirectional transmission with $\mathbf{R}_{X}=\mathbf{I}$ to calculate the fixed receive beamformer $\mathbf{w}$ according to (16) in the simulations. Once the optimal transmit covariance matrix $\mathbf{R}_{X}^{*}$ is obtained, we substitute it into (16) to derive the corresponding optimal receive beamformer $\mathbf{w}^{*}$.
@@
We investigate the impact of different subarray distributions on the performance of ISAC systems. To evaluate the impact of array geometry, we compare three modular array configurations under fixed $K$ and $M$: random layout where subarrays are randomly distributed, uniform layout where subarrays are uniformly distributed, and collocated layout where subarrays are closely spaced with half-wavelength separation. Fig. 4 shows the communication spectral efficiency versus received SNR for proposed algorithms under different array configurations. It is noted that for both algorithms, the modular arrays with random and uniform layouts achieve similar performance and significantly outperform the collocated layout, with the performance gap widening at higher SNRs. This superiority of widely-spaced distributions stems from their enhanced near-field characteristics, leading to higher channel rank and improved spatial multiplexing gain. Additionally, SDR-RRS consistently outperforms RM-JGD across all configurations, demonstrating its better capability in handling the non-convex optimization problem.
Figure 5: Communication spectral efficiency versus sensing SCNR threshold for different beamforming schemes.
We compare the proposed SDR-RRS and RM-JGD algorithms with three baselines: fully-digital beamforming (FDB) as the performance upper bound, subarray-based alternating algorithm (SAO) [17], and consensus alternating direction method of multipliers (CADMM) [33]. Fig. 5 shows the communication spectral efficiency versus sensing SCNR threshold for different schemes with 28 RF chains. The results reveal that SDR-RRS consistently achieves the highest spectral efficiency among all hybrid beamforming schemes, closely approaching the FDB upper bound. Additionally, CADMM outperforms RM-JGD at low SCNR thresholds, while SAO shows the lowest performance. The SDR-RRS algorithm’s superior performance stems from its ability to find a near-optimal solution by leveraging the optimal analog beamformer and approximating the global optimum, while the RM-JGD algorithm may converge to a local optimum far from the global optimum, resulting in slightly inferior performance. In contrast, the SAO and CADMM algorithms alternately optimize the digital and analog beamformers and employ various approximations to the objective function and rely on various approximations, which may lead to suboptimal solutions and further performance degradation.
@@
Fig. 7 investigates the impact of subarray scale on the communication performance of SDR-RRS algorithm. Under fixed total array aperture $S$ and antenna number $N_{t}=192$, we vary the subarray antenna number $M$ as 16, 24, 32, 48, and 64, with corresponding subarray number $K$ being 12, 8, 6, 4, and 3. It can be observed that the optimal subarray scale varies with the user distance. For users within 25 m, the subarray scale of $M=48$ outperforms the others, while beyond 25 m, the subarray scale of $M=64$ achieves the highest communication spectral efficiency. A larger number of subarrays $K$ provides higher spatial multiplexing gains, but a smaller $M$ reduces the beamforming gain. Thus, the optimal subarray scale at different distances represents a trade-off between spatial multiplexing and beamforming gains for maximizing the communication performance. Moreover, the performance gap between subarray scales decreases with distance, due to the transition from near-field spherical wavefronts to far-field planar wavefronts, which gradually diminishes the spatial multiplexing advantages of having more subarrays.
In Fig. 8, the normalized MUSIC spectrum obtained by the proposed SDR-RRS algorithm is compared for $M=32$ with varying numbers of subarrays $K$, over a grid spanning $x\in[0:0.06:30]$ m and $y\in[0:0.06:30]$ m. The peak of the spectrum is consistently observed at the actual target location $(20\mathrm{m},\pi/4)$ for all values of $K$. Notably, the main lobe width containing the peak value narrows as $K$ increases, indicating enhanced range resolution at the same angular direction. This behavior can be attributed to the spherical wavefront characteristic across subarrays, which becomes more pronounced with increasing $K$, leading to improved range estimation accuracy. The results demonstrate that augmenting the number of subarrays can significantly enhance the range resolution capability of the MUSIC algorithm, thereby enabling more precise target localization.
@@ VI Conclusion
In this paper, we have developed the low-complexity hybrid beamforming algorithms for modular XL-MIMO ISAC systems. By exploiting the structural similarity between the communication and sensing channels based on the the piecewise-far-field channel model, we have derived a closed-form solution for the optimal analog beamformer and transformed the joint analog-digital design problem into a lower-dimensional digital beamformer optimization. Two approaches have been developed for the rank-constrained digital beamformer optimization: a manifold-based method and an SDR-based method for obtaining near-optimal solutions. Extensive simulations have validated the effectiveness of the proposed algorithms.
Appendix A Proof of Lemma 1
First, since $\mathbf{H}_{c}\in\mathbb{C}^{N_{c}\times N}$ and $N_{c}\ll N$, the rank of $\mathbf{H}_{c}$ is upper bounded by $N_{c}$. For a given propagation path $p$, the subarray response vectors $\mathbf{a}_{tp}^{1}\left(\theta_{tp}^{1}\right),\ldots,\mathbf{a}_{tp}^{K}\left(\theta_{tp}^{K}\right)$ are linearly independent. Similarly, due to the distinguishability of $N_{p}$ propagation paths, the angles of different paths are different. Therefore, for a fixed transmit subarray $k$, the response vectors $\mathbf{a}_{t1}^{k}\left(\theta_{t1}^{k}\right),\ldots,\mathbf{a}_{tN_{p}}^{k}(\theta_{tN_{p}}^{k})$ for different propagation paths $p=1,\ldots,N_{p}$ are linearly independent. Therefore, we can obtain the $n$-th row of the communication channel $\mathbf{H}_{c}$ in (5) $\mathbf{H}_{c}(n,:)$ is given by
${{\mathbf{H}}_{c}}\left(n,:\right)=$ (55)
$\left[\sum\limits_{p=1}^{{{N}_{p}}}{\mu_{p}^{1}}\mathbf{a}_{cp}^{1}\left(n\right){{\left(\mathbf{a}_{tp}^{1}\right)}^{H}}\right.\left.\cdots\ \sum\limits_{p=1}^{{{N}_{p}}}{\mu_{p}^{K}}\mathbf{a}_{cp}^{K}\left(n\right){{\left(\mathbf{a}_{tp}^{K}\right)}^{H}}\right],$
where $\mathbf{a}_{tp}^{k}$ is a shorthand notation for $\mathbf{a}_{tp}^{k}\left(\theta_{tp}^{k}\right)$, and $\mathbf{a}_{cp}^{k}\left(n\right)$ denotes the $n$-th element of $\mathbf{a}_{cp}^{k}\left(\theta_{cp}^{k}\right)$. Therefore, each row vector $\mathbf{H}_{c}(n,:)$ is a linear combination of the following $KN_{p}$ linearly independent row vectors:
$\mathbf{\bar{a}}_{tp}^{k}\triangleq[\ \underbrace{\mathbf{0},\ldots,\mathbf{0}}_{k-1},(\mathbf{a}_{tp}^{k})^{H},\underbrace{\mathbf{0},\ldots,\mathbf{0}}_{K-k}\ ]\in\mathbb{C}^{1\times KM},$ (56)
where $k=1,2,\cdots,K$, $p=1,2,\cdots,N_{p}$, and $\mathbf{0}$ is an all-zero vector of dimension $1\times M$.
However, the arrival angles corresponding to different propagation paths from different transmit subarrays may not be spatially distinguishable at the receive array, resulting in potential linear dependency among the vectors in (56). This dependency reduces the rank of the channel matrix $\mathbf{H}_{c}$, thereby establishing its upper bound as $KN_{p}$, which validates the right-hand inequality. The left-hand inequality holds in the extreme case where all transmit subarrays share the same arrival angle to the receive array for a given propagation path $p$, i.e., $\theta_{cp}^{1}=\theta_{cp}^{2}=\ldots=\theta_{cp}^{K}$[34].
Appendix B Proof of Theorem 1
By letting $\mathbf{R}_{X}^{*}=\bm{\Delta}\bm{\Delta}^{\mathrm{H}}$, we can decompose $\bm{\Delta}$ as